Fermi National Laboratory
 

Operation Summary for February 23, 2011 - February 25, 2011

 

Wednesday Feb. 23

The day shift began with Operations monitoring the Tevatron (TEV) store 8526 (80.51E30) and the Recycler stash (247.2E10), with the Antiproton Source (Pbar) stacking (26.3E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Operations terminated store 8526 at 2:30 PM. Operations established store 8527 at 3:50 PM with an average initial luminosity (AIL) of 334.2E30. Operations halted the MI beam for twenty minutes to reboot MECAR. Problems with a Linac RF station (LRF4) held beam off for an hour.

During the evening shift, Operations monitored the accelerator complex.

Thursday Feb. 24

The midnight shift began with Operations monitoring store 8527 (118.28E30), the stash (207.3E10), and the Pbar stack (23.5E10), with NuMI off due to a target LCW leak, and with MiniBooNE taking beam.

During the midnight shift, at 4:30 AM, a cryo system failure for the Pbar stochastic cooling system halted stacking. At 4:45 AM, NuMI tripped off due to a problem with its target LCW system. At 7:55 AM, a transformer failure in the MI-8 service building halted beam to MiniBooNE and aborted the Recycler stash of 311E10. Pbar and NuMI remained off for the rest of the midnight shift.

During the day shift, operators worked at recovering from the MI-8 power glitch. Operations restored beam to MiniBooNE at 11:43 AM. Operators reported that MI personnel accessed the tunnel to repair a faulty beam valve limit switch. Pbar remained off for cryo system repairs. MiniBooNE took beam for about half an hour.

During the evening shift, operators had MI on by 5 PM. NuMI personnel performed target scans until 6 PM. MI and Controls personnel addressed a series of MECAR (MI main bus regulation) problems. Operations prepared NuMI for an access to investigate the target LCW leak. Pbar was ready for beam by 10 PM, but unfortunately, more MECAR related issues held off MI for the remainder of the shift. MiniBooNE took beam for about 7.5 hours.

Friday Feb. 25

The midnight shift began with Operations monitoring store 8527 (37.92E30), with no stashing or stacking due to MECAR problems, with NuMI off due to a target LCW leak, and with MiniBooNE taking beam.

During the midnight shift, Operations reported that MI and Controls personnel resolved the MECAR problems and Operations resumed sending beam to MI at 1:50 AM. MI beam reached its normal intensity at 3:30 AM. Operations reported that a Pbar Debuncher injection power supply problem held off stacking until 6 AM. MiniBooNE ran all shift.

The Plans for Friday and the Weekend

The plans for today are to resume normal stacking, stashing, and storing. The Run Coordinator will allow TEV personnel to conduct end-of-store studies.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/25/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory